Two angles are complementary.One angle is six less than twice the measure of the other. What is the measure of each angle.
Answer by nerdybill(7003)

Two angles are complementary.One angle is six less than twice the measure of the other. What is the measure of each angle.
.
Let x = angle
then
90-x = complementary of angle
.
x = 2(90-x) - 6
x = 180-2x - 6
x = 174-2x
3x = 174
x = 58 degrees
.
complementary:
90-x = 90-58 = 32 degrees
.
answer: 32 and 58 degrees